Write first five multiples of: 
8 ??????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The required multiples are:
8 ×1 =8
8 × 2 = 16
8 × 3 = 24
8 × 4 = 32
8 × 5 = 40
Hence, the first five multiples of 8 are 8, 16, 24, 32 and 40

Thanks